OpenStudy (anonymous):

i need help what the real solution to 6x=9+x^2

OpenStudy (anonymous):

All you need to do is simplify. 6x-x^(2)=9 -x^(2)+6x=9 -x^(2)+6x-9=0 x^(2)-6x+9=0 (x-3)(x-3)=0 (x-3)^(2)=0 x-3=0 x=3

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

No problem c: